You are here: Foswiki>PartnerObserving Web>ItalyQueue2012>ObsInaf20121209 (28 Feb 2013, EleanoraSani)Edit Attach

INAF LBC+MODS+LUCI queue observing, 2012 December 8-9

Observer: E. Sani, F. Cusano, M. Fumana
Telescope Operator: G. Becchetti
Support Astronomer: O. Kuhn
Instrument: LUCI/MODS

Summary

Before sunset there are cirrus and clouds approaching from west/sauth-west.
During the twilight D. Thomson tests the LUCI stability.
At the beginning of the night we have seeing 1" and 33% humidity with low wind.
LUCI looks to have a systematic offset and MOS ACQU could not works properly. S.A. suggests us to correct this error applying a x=0.5 offset and PA=0.15 rotation. It is not clear if this solution works. This is because we actually do not know what MOS ACQU actually does: the rotation center is unknown.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Telluric Hipp107734 K, 210 zJHK    
Cresci_A DSF2237a K, 210 zJHK 1h 10min COMPLETED
Pentericci_A CLUSTERPENTE HKspec, 200_H+K  
Fritz_04A LensGG243 dual grating 2h25min COMPLETED
Standard G191-B2B red grating    
Standard feige34 red grating   cirrus passing
Gilli_B       Aborted

00.37 The T.O. opens the Dome

01.35 executeLuciscript.sh clean1.txt
  • We flush the detector
  • We test again LUCI checking the background counts in a dark frame

01:47 executeLuciscript.sh tell2_acq_campoA.txt
  • error message from filter whell #2
  • press INITIALIZE on the Instrument Control Panel. Failed. S.A. fix the problem

01.59 executeLuciscript.sh tell2_acq_campoA.txt luci.20121209.0006-0009

02.08 executeLuciscript.sh tell_sci_campoA_K.txt luci.20121209.0010-0023
  • wrong rotation, we have to abort the acquisition

02.51 executeLuciscript.sh tell_sci_campoA_K.txt luci.20121209.0024-0028

03.07 executeLuciscript.sh sci_campoA_K_1.txt luci.20121209.0029-000040

04.15 executeLuciscript.sh sci_campoA_K_1.txt luci.20121209.0041-0042

04.27 We slew to Pentericci_02
  • executeLuciscript.sh pente1.acq luci.20121209.0043-0046
  • the average seeing is 0.6"

04.47 executeLuciscript.sh pente1_HK_1.spec luci.20121209.0047
  • Aborted. We try to give another offset to move object in slit #9
  • Get an acquisition image manually luci.20121209.0048

05:05 executeLuciscript.sh pente1_HK_1.spec luci.20121209.0049-0053

05.35 PROBLEM Loose of the guidestar. Restarting GCS.

06:02 executeLuciscript.sh pente1.acq. FAILED preset cancel

06:03 executeLuciscript.sh pente1.acq luci.20121209.0049-0054
  • Problem in computation of offset and PA rotation. Try to perform compute by hand

06:30 executeLuciscript.sh pente1_HK_1.spec luci.20121209.0058-0061
  • Airmass ~1.6". Script interrupted

06:54 SWITCH TO MODS
  • GCS left telemetry disconnected. The MODS calibration unit was in.

Problems with the pointing. It seems that the MODS guide probe is not in the correct position

08.04 We are authorized to use MODS and slew to Fritz_04A.

08.17 execMods FritzLensGG243.obs ==mods1r.20121209.0008-0010 / mods1b.20121209.0002-0004
  • the averagwe seeing is 1.0"

10.42 We slew to the standard star of Gilli_A
  • acqMods standard.for.cxoj09.g191b2b.acq ==mods1r.20121209.0012-0013
  • modsAlign -r mods1r.20121209.0012

10.51 execMods standard.for.cxoj09.g191b2b.obs ==mods1r.20121209.0014-0016

11.01 We slew to the scientific target of Gilli_A *acqMods cxoj09.acq ==mods1r.20121209.0016-0019 target to faint to be seen on the acquisition image, we double the exposure time, we still don't see the target

11.18 We slew to the scientific target of Gilli_B
  • acqMods 2xmmj10.acq ==mods1r.20121209.0020-0022
  • In the acquisition image we don't see the target, but we identified the field and we believe the the target is in the slit
  • pointing problem we have to launch again the acquisition

11:45 acqMods 2xmmj10.acq mods1r.20121209.0023
  • Preset cancel. Telescope problem. Run acq again
  • seeing 1.3-1.6 we are have problem to collimate the telescope

12:05 execMods 2xmmj10.obs ==mods1r.20121209.0024

12:25 No Object appears in slit. Abort

12:27 acqMods feige34.acq mods1r.20121209.0025-0026
  • modsAlign -r

12:36 execMods feige34.obs mods1r.20121209.0028-0029
  • we have some cirrus passing for this standard

13:00 Dome close

13:01 Start calibration mods1r.20121209.0030-00??
  • execMods grpixflats.cal
  • execMods grpixflats2.cal
  • execMods grmosflats.cal
  • execMods grmoslamps.cal
  • execMods grflats50.cal. Satureted frames: mods1r.20121209.0048-mods1b.20121209.0023
  • execMods grflats10.cal
  • execMods grlamps10.cal

14.50 execMods modsSlip.pro
  • execMods bias8K.cal

-- EleanoraSani - 08 Dec 2012
Topic revision: r5 - 28 Feb 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback